OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP!! WILL MEDAL AND FAN!! Which expression is equivalent to (cos x)(tan(-x))? -sin x sin x -csc x csc x

OpenStudy (anonymous):

@stgzlg could you please help?

OpenStudy (anonymous):

\[\cos x=\cos \left( -x \right)\]

OpenStudy (anonymous):

\[\tan \left( -x \right)=\frac{ \sin \left( -x \right) }{ \cos \left( -x \right) }\] \[\sin \left( -x \right)=-\sin x\]